OpenStudy (a1234):

I'm asked to show the calculations proving momentum is conserved with this data for a nonlinear collision.

OpenStudy (a1234):

What do I do with the Vx and Vy?

OpenStudy (3mar):

@welshfella Call of duty here

OpenStudy (welshfella):

Momentum = mass * velocity

OpenStudy (welshfella):

the total momentum before the collision should equal the total momentum after the collision.

OpenStudy (welshfella):

Momentum before the collision = 0.500 * 0.700 = 0.35

OpenStudy (a1234):

What about the y-component?

OpenStudy (welshfella):

the y component is 0

OpenStudy (welshfella):

after collision x component = 1.5 * 0.341 + (0.5 -0.322) y component = 0.5 * 0.170 + (1.5 8 -0.057) the resulting momentum should equal 0.35

OpenStudy (a1234):

The x-component is 0.3505 and the y-component is -0.0005. Are these supposed to be added?

OpenStudy (welshfella):

yes

OpenStudy (a1234):

So then we get 0.35

OpenStudy (welshfella):

- which gives you 0.35 for the after collision total momentum

OpenStudy (welshfella):

yea

OpenStudy (a1234):

The momentum before and after the collision are equal, so it is conserved

OpenStudy (welshfella):

momentum is always conserved.

OpenStudy (welshfella):

right

OpenStudy (a1234):

Thanks!

OpenStudy (welshfella):

yw note that momentum is a vector so its direction has to be taken into account momentum to the right is usually taken to be positive.

OpenStudy (a1234):

Yes, the positive direction is indicated on the pictures above

OpenStudy (welshfella):

so they are.

OpenStudy (3mar):

Thank you for good explanation. @welshfella